The vapor pressure of water above a solution of water and a nonvolatile solute at 25.0 °C is 21.0 mm Hg. What is the mole fraction of the solute? (The vapor pressure of pure water is 23.8 torr at 25.0 °C).

1
Expert's answer
2021-02-05T04:04:08-0500

χs = Δp / p0 = (23.8-21.0)/23.8 = 0.1176